### v1

# AltFS: Agency-light Feature Selection with Large Language Models in Deep Recommender Systems

Abstract
Feature selection is crucial in recommender systems for improving model efficiency and predictive performance. Traditional methods rely on agency models, such as decision trees or neural networks, to estimate feature importance. However, this approach is inherently limited, as the agency models may fail to learn effectively in all scenarios due to suboptimal training conditions (e.g., feature collinearity, high-dimensional sparsity, and data insufficiency). In this paper, we propose AltFS, an Agency-light Feature Selection method for deep recommender systems. AltFS integrates semantic reasoning from Large Language Models (LLMs) with task-specific learning from agency models. Initially, LLMs will generate a semantic ranking of feature importance, which is then refined by an agency model, combining world knowledge with task-specific insights. Extensive experiments on three public datasets from real-world recommender platforms demonstrate the effectiveness of AltFS. Our code is publicly available  [^1] The link will be released after publication for reproducibility.
AltFS: Agency-light Feature Selection with Large Language Models in Deep Recommender Systems
Pengyue Jia1, Zhaocheng Du2, Yichao Wang2, Xiangyu Zhao1, Xiaopeng Li1, Yuhao Wang1, Qidong Liu1, Huifeng Guo2, Ruiming Tang2, 1City University of Hong Kong, 2Huawei Noah’s Ark Lab
1 Introduction
Feature selection is crucial in deep recommender systems (DRS) for improving model performance, mitigating overfitting, and accelerating training and inference Wang et al. (2022); Lin et al. (2022). Numerous solutions have been proposed to address the need for feature selection in DRS, which can be broadly categorized into three types based on their selection techniques: (1) Shallow feature selection methods Breiman (2001); Chen and Guestrin (2016), which typically use statistical algorithms to assign importance scores to features. (2) Gate-based feature selection methods Guo et al. (2022); Lyu et al. (2023), where a gate vector is optimized during training by assigning gates to feature embeddings, and the gate values are interpreted as indicators of the corresponding features’ importance. (3) Sensitivity-based feature selection methods Wang et al. (2023); Zhang et al. (2023); Du et al. (2024), which evaluate parameter sensitivity using gradient information during backward propagation and then calculate feature importance.
Figure 1: Conventional paradigm and AltFS.
As shown in Figure 1, all the aforementioned methods rely on training an agency model to capture the feature-to-label relationship. The accuracy of feature selection, therefore, hinges on how closely the agency model’s mapping aligns with the ground truth. However, in many real-world recommendation scenarios, agency models often fail to achieve optimal performance. For example, in cold-start or deep conversion tasks, the sparsity of samples—particularly positive samples—frequently leads to underfitting Schein et al. (2002). Conversely, in cases involving numerous high-cardinality features, agency models are prone to overfitting Zhang et al. (2022). Moreover, these models typically struggle to account for feature interdependencies, overlooking critical factors such as feature collinearity and complementarity when estimating feature importance.
Figure 2: Overview of AltFS. There are three stages in AltFS, (1) Feature Importance Extraction Stage, (2) Feature Importance Refinement Stage, and (3) Retraining Stage.
LLMs offer potential solutions to the challenges faced by agency models. Trained on vast amounts of web data, LLMs can leverage their world knowledge to identify informative feature subsets in sparse scenarios and comprehend feature dependencies based on feature descriptions. For instance, agency models may fail to recognize that “longitude” and “latitude” must coexist to uniquely identify positions, potentially selecting only one for the feature subset. However, when this understanding is integrated into an LLM’s world knowledge, such issues can be easily resolved.
However, several challenges remain: (1) Feature Complexity: In real-world online platforms, numerous features exhibit complex relationships both among themselves and with the tasks at hand. (2) Knowledge Gap: LLMs are generally pre-trained on broad world knowledge, which may not align with the specific requirements of recommendation tasks. (3) Efficiency Demand: Additionally, industrial scenarios impose strict requirements on efficiency and resource consumption. Thus, efficiently and lightweightly integrating feature importance from world knowledge into recommendation model training poses a significant challenge.
To address the challenges above, we propose AltFS, an Agency-light Feature Selection algorithm for DRS. Specifically, to explore complex feature interdependencies and feature-task relationships, we design a context-aware prompt iteration method that leverages LLMs to provide prior knowledge of feature importance. To efficiently and lightweightly integrate this knowledge into agency models for assessing feature importance, we introduce a bridge network to refine the feature importance rankings further. The bridge vector is trained under constraints to harmonize knowledge-based and task-specific feature importance. Our contributions can be summarized as follows:
  • AltFS introduces the first paradigm that combines world knowledge priors with agency-based feature selection in DRS, effectively addressing traditional methods inaccuracies that arise from relying solely on the agency model.
  • We design a prompt iteration method that enables LLMs to select effective features based on their world knowledge iteratively.
  • We introduce a novel bridge network to integrate feature importance from world knowledge and the recommendation task space. This network is optimized in a lightweight, end-to-end manner.
  • Extensive experiments are conducted on three public datasets derived from real-world online service platforms, verifying the effectiveness of our method.
2 Methodology
In this section, we will detail AltFS. First, we provide an overview of the AltFS framework in Section 2.1. Next, we introduce the fundamental DRS model in Section 2.2 and the key modules of AltFS in Section 2.3 and Section 2.4. Finally, we will describe the optimization and retraining in Section 2.5 and Section 2.6.
2.1 Overview
In this subsection, we provide an overview of AltFS, which consists of three stages: (1) Feature Importance Extraction, (2) Feature Importance Refinement, and (3) Retraining, as shown in Figure 2. In the Feature Importance Extraction stage, a prompt iteration method leverages LLMs to iteratively select predictive features from the candidate set, producing an initial feature importance ranking. To reduce bias, multiple LLMs are employed concurrently. Next, in the Feature Importance Refinement stage, the extracted feature importance sequences are refined using a bridge network integrated into the training process of the recommendation model. Finally, in the Retraining stage, the refined feature importance sequences are combined using the optimized bridge vector, generating a ranking that incorporates semantic reasoning and aligns with the recommendation task. The top $n$ features from this ranking are selected to retrain a recommendation model from scratch.
2.2 Basic DRS Architecture
2.2.1 Embedding Layer
Efficient handling of categorical data is vital in recommender systems due to its inherent sparsity. Embedding layers address this challenge by transforming high-dimensional sparse categorical inputs into dense, low-dimensional vectors for DRS. The embedding process consists of two steps: (1) Binarization: Categorical features $\boldsymbol{x}$ are converted into binary vectors $\boldsymbol{x}^{\prime}$ using one-hot encoding, where the dimensionality depends on the number of unique values in each feature field. (2) Projection: The binary vectors are projected into dense embeddings via embedding table lookups. For each feature field, the binary vector $\boldsymbol{x}_{n}^{\prime}$ is mapped to an embedding $\boldsymbol{e}_{n}=\boldsymbol{x}_{n}^{\prime}\cdot\boldsymbol{M}_{n}$, where $\boldsymbol{M}_{n}$ is the learnable embedding table, and $\boldsymbol{e}_{n}\in\mathbb{R}^{dim}$ is the resulting embedding. The concatenated output is $\boldsymbol{E}=[\boldsymbol{e}_{1},\boldsymbol{e}_{2},\cdots,\boldsymbol{e}_{N}]$, where $\boldsymbol{E}$ represents the combined feature embeddings.
2.2.2 Transformation Layer
The transformation layer is a key component in DRS, enabling robust fitting capabilities through linear transformations and nonlinear activation:
$\boldsymbol{h}_{l}=\text{ReLU}(\boldsymbol{W}_{l-1}\boldsymbol{h}_{l-1}+\boldsymbol{b}_{l-1}),\ \boldsymbol{h}_{0}=\boldsymbol{E}$ (1)
$\hat{y}=\sigma(\boldsymbol{W}_{L}\boldsymbol{h}_{L}+\boldsymbol{b}_{L})$ (2)
where $\boldsymbol{h}_{l}$ is the hidden state of the $l$th layer, and $\boldsymbol{W}_{l}$ and $\boldsymbol{b}_{l}$ are the weight and bias for the $l$th layer. For the output layer, $\boldsymbol{h}_{L}$, $\boldsymbol{W}_{L}$, and $\boldsymbol{b}_{L}$ denote the hidden state, weight, and bias, respectively. $\sigma$ represents the activation function for recommendation tasks (e.g., Sigmoid and Softmax).
2.3 Prompt Iteration
To extract the complex semantic information between features and their relevance to specific recommendation tasks, we design the prompt illustrated in Figure 3, which consists of five parts: instructions, descriptions, feature sets, supplementary information, and output formatting. (1) Instructions: This section defines the role of LLMs and introduces the subsequent tasks. (2) Descriptions: This part provides essential information to address the feature selection problem, including descriptions of the recommendation task, dataset, and feature set. Specifically, in the recommendation task description, we outline the background of the task. For example, in the MovieLens dataset, the task is to determine whether a user likes a movie based on user profiles and movie information. The dataset description focuses on the number of samples and features. Finally, the feature set description is divided into user features, item features, and interaction features. Each feature description includes five elements: feature name, feature description, feature type (e.g., int, string), example, and unique values, providing sufficient information for LLMs to assess feature importance. (3) Feature Sets: The feature sets are composed of two parts: features already selected and candidate features. The union of these two sets constitutes the complete feature set. (4) Supplementary Information: This part emphasizes the expected actions of LLMs (selecting one feature from the candidate features to add to the already selected features during each inference) and provides reference criteria for feature selection. (5) Output Formatting: This section standardizes the output format of LLMs to facilitate automated and streamlined feature selection. The prompts we used in this work are detailed in Appendix A.1.
Figure 3: Structure of prompt.
Inspired by existing work Sun et al. (2023), we prompt LLMs to analyze and select informative features in an iterative manner. Considering that different LLMs may exhibit significant assessing bias due to variations in training data, we treat the LLM as an expert while simultaneously introducing multiple different LLMs to generate several feature importance sequences:
$f_{t}^{k}=\text{LLM}_{k}(\text{prompt},\boldsymbol{s}_{t}^{k},\boldsymbol{c}_{t}^{k})$ (3)
$\boldsymbol{s}_{t+1}^{k}=\boldsymbol{s}_{t}^{k}\cup\{f_{t}^{k}\},\ \boldsymbol{c}_{t+1}^{k}=\boldsymbol{c}_{t}^{k}\setminus\{f_{t}^{k}\}$ (4)
where $\boldsymbol{s}_{t}^{k}$ and $\boldsymbol{c}_{t}^{k}$ denote the set of selected features and the set of candidate features at $t$th step with $k$th LLM, $f_{t}$ is the selected feature at $t$th step with $k$th LLM. The feature selected at $t$th step will be used to update the set of selected features and candidate features through Equation 4. Since user ID and item ID are essential in recommendation tasks, we initialize $\boldsymbol{s}_{0}$ as a set containing these two features. The iteration process continues until no features remain in the candidate feature set. The output is a matrix containing the features selected at each time step by different LLMs.
2.4 Bridge Network
The feature importance ranking obtained through prompting LLMs is based solely on general world knowledge, leading to a discrepancy with the feature importance derived during model training. To bridge this gap, we propose a bridge network that incorporates feature importance sequences produced by different LLMs into the training process of the recommendation model.
For batch data, a few features will be first masked according to the feature importance ranking produced by LLMs. The number of masked features $r$ varies in each batch and is sampled from $r\sim\mathcal{U}(0,N\cdot\beta)$, where $N$ and $\beta$ are the number of feature fields and maximum mask ratio.
[…]
where $\boldsymbol{\mathcal{M}}$ is the masking matrix with $m_{t}^{k}$ representing the masking value at $t$th step for $k$th LLM.
To integrate feature importance rankings produced LLMs into recommendation model training, we introduce a bridge vector. Suppose the feature set is $F$, $\boldsymbol{w}$ is the learnable bridge vector $\boldsymbol{w}=[w_{0},w_{1},\cdots,w_{K}]$ with a length equal to the number of LLMs, the integration process is:
$\delta(\boldsymbol{F}_{n},f_{t}^{k})=\begin{cases}1&\text{if}\ \boldsymbol{F}_{n}=f_{t}^{k}\\ 0&\text{otherwise}\end{cases}$ (6)
$\boldsymbol{\hat{w}}=\text{Softmax}(\boldsymbol{w}\cdot\text{exp}(\tau))$ (7)
$\boldsymbol{h}_{n}=\sum\nolimits_{k=0}^{K}\sum\nolimits_{t=0}^{N}\delta(\boldsymbol{F}_{n},f_{t}^{k})\cdot(m_{t}^{k}\cdot\boldsymbol{\hat{w}}_{k})$ (8)
where $\boldsymbol{F}_{n}$ is the $n$th feature field. $\boldsymbol{h}_{n}$ denotes the integrated weight for $n$th feature field. $K$ and $N$ denote the number of LLMs and feature fields, and $\tau$ is the temperature.
Before assigning the integrated weights $\boldsymbol{h}$ to embeddings, we need to normalize embeddings because their magnitude varies significantly across different feature fields, which can not be compared directly. We conduct BatchNorm to each field embedding by $\hat{\boldsymbol{e}}_{n}=(\boldsymbol{e}_{n}-\text{E}_{n})/\sqrt{\text{V}_{n}+\epsilon}$, where $\text{E}_{n}$ and $\text{V}_{n}$ are the mean and variance of $n$th feature field in a batch, $\epsilon$ is a constant introduces to enhance numerical stability. After transformation, the embeddings of features should be $\boldsymbol{\hat{E}}=[\boldsymbol{\hat{e}}_{1},\boldsymbol{\hat{e}}_{2},\cdots,\boldsymbol{\hat{e}}_{N}]$. Next, we weights embeddings by $\boldsymbol{E}^{\prime}=[\boldsymbol{h}_{1}\boldsymbol{\hat{e}}_{1},\boldsymbol{h}_{2}\boldsymbol{\hat{e}}_{2},\cdots,\boldsymbol{h}_{N}\boldsymbol{\hat{e}}_{N}]$, where $\boldsymbol{E}^{\prime}$ denotes weighted embeddings, $\boldsymbol{h}_{n}$ and $\boldsymbol{\hat{e}}_{n}$ are the weight and embeddings after transformation for $n$th feature field. The weighted embeddings will be input into the transformation layer in DRS to generate predictions following Equation 1 and Equation 2.
2.5 Optimization
We take the Click-through rate (CTR) prediction as our task in this paper, and we use the binary cross-entropy (BCE) loss as our optimization objective:
$\text{min}\ \mathcal{L}=-(\boldsymbol{y}_{i}\text{log}(\hat{\boldsymbol{y}}_{i})+(1-\boldsymbol{y}_{i})\text{log}(1-\hat{\boldsymbol{y}}_{i}))$ (9)
where $\boldsymbol{y}$ and $\hat{\boldsymbol{y}}$ are the ground truth label and the prediction value.
2.6 Retraining
Given the well-trained bridge vector $\boldsymbol{w}^{*}$ and weighting gate $\boldsymbol{\hat{w}}^{*}$, we can obtain the final feature importance scores for each feature, which simultaneously consider general world knowledge and recommendation task information.
$\boldsymbol{fi}_{t}^{k}=1-\frac{t}{N}$ (10)
$\boldsymbol{h}_{n}^{*}=\sum\nolimits_{k=0}^{K}\sum\nolimits_{t=0}^{N}\delta(\boldsymbol{F}_{n},f_{t}^{k})\cdot(\boldsymbol{fi}_{t}^{k}\cdot\boldsymbol{\hat{w}}_{k}^{*})$ (11)
where $\boldsymbol{fi}_{t}^{k}$ denotes the feature importance score for feature $\boldsymbol{f}_{t}^{k}$ with $k$th LLM, the score decreases linearly according to the feature’s ranking in the sequence. $\boldsymbol{h}_{n}^{*}$ and $\boldsymbol{\hat{w}}_{k}^{*}$ are the final integrated feature importance score for $n$th feature field and well-trained weighting gate for $k$th LLM expert.
The top-$n$ feature fields with the highest $\boldsymbol{h}_{n}^{*}$ will be selected to retrain DRS from scratch.
Table 1: Comparison results between AltFS and other state-of-the-art baselines on three public datasets. The metrics of the best-performed methods and sub-optimal methods are highlighted in bold fonts and underlined. “*” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. $\uparrow$: higher is better. $\downarrow$: lower is better.
Dataset Movielens-1M Aliccp Kuairand
AUC $\uparrow$ Logloss $\downarrow$ AUC $\uparrow$ Logloss $\downarrow$ AUC $\uparrow$ Logloss $\downarrow$
No Selection 0.78854 0.54367 0.61804 0.16190 0.77953 0.55872
Shallow Feature Selection
Lasso 0.78853 0.54388 0.61861 0.16185 0.77987 0.55835
GBDT 0.78904 0.54318 0.61854 0.16232 0.77972 0.55835
RF 0.78868 0.54336 0.61910 0.16172 0.77988 0.55801
XGBoost 0.80081 0.53075 0.61873 0.16176 0.77992 0.55806
Gate-based Feature Selection
AutoField 0.80315 0.53075 0.61922 0.16141 0.77997 0.55820
AdaFS 0.78523 0.54765 0.61869 0.16143 0.77981 0.55854
OptFS 0.77606 0.55564 0.61862 0.16264 0.77933 0.55855
LPFS 0.80339 0.52822 0.61836 0.16189 0.77983 0.55824
Sensitivity-based Feature Selection
Permutation 0.78853 0.54354 0.61935 0.16152 0.77978 0.55831
SFS 0.78927 0.54241 0.61897 0.16180 0.77981 0.55819
SHARK 0.78904 0.54277 0.61928 0.16151 0.77995 0.55815
AltFS 0.80480 0.52703 0.62015 0.16139 0.78002 0.55801
3 Experiments
3.1 Experimental Settings
3.1.1 Dataset.
We conduct experiments on three public datasets: Movielens-1M [^2] https://grouplens.org/datasets/movielens/1m/, Aliccp Ma et al. (2018), and Kuairand Gao et al. (2022). These datasets have varying numbers of features (9 in Movielelns-1M, 23 in Aliccp, and 102 in Kuairand). They are chosen to illustrate the effectiveness of our method under different application conditions. Detailed descriptions of datasets are given in Appendix A.2.
3.1.2 Baselines.
We perform experiments with the following baseline methods to demonstrate the advanced performance of our method. (1) Shallow feature selection methods: Lasso Tibshirani (1996), GBDT Friedman (2001), RandomForest Breiman (2001), XGBoost Chen and Guestrin (2016). (2) Gate-based feature selection methods: AutoField Wang et al. (2022), AdaFS Lin et al. (2022), OptFS Lyu et al. (2023), LPFS Guo et al. (2022). (3) Sensitivity-based feature selection methods: Permutation Fisher et al. (2019), SFS Wang et al. (2023), SHARK Zhang et al. (2023). The detailed information of these baselines is introduced in Appendix A.3.
3.1.3 Evaluation Metrics.
To comprehensively evaluate our method, we use the AUC and Logloss metrics in CTR prediction following previous work Wang et al. (2022); Lin et al. (2022). It is worth noting that an increase of 0.001 in AUC or a decrease in Logloss is already a significant improvement in CTR prediction task Song et al. (2019); Wang et al. (2021).
3.1.4 Implementation Details.
We use Adam optimizer with learning rate 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=1\times 10^{-8}$ for all experiments. The batch size is 4096, the embedding size is 8, the number of LLMs $K$ is 3, the maximum mask ratio is 0.2, and the temperature coefficient $\tau$ is 4.0. To minimize the impact of randomness, we conduct each trial three times and use the average as the final result. The detailed settings are introduced in Appendix A.4.
3.2 Overall Performance
In this section, we compare AltFS with other state-of-the-art baselines on three public datasets. The results are listed in Table 1, and we can derive the following findings: 1) AltFS achieves the best performance across all metrics on three datasets. This superior performance is attributed to AltFS’s unique approach, which does not rely solely on the information provided by agency models to determine feature importance. Instead, it combines both world knowledge and task-specific information to comprehensively assess feature importance. 2) Overall, Gate-based and Sensitivity-based feature selection methods outperform Shallow feature selection methods. This is because Shallow feature selection methods rely on the original classifier, whose performance is generally weaker compared to the deep neural networks employed by Gate-based and Sensitivity-based feature selection methods. 3) The performance improvement of AltFS compared to no selection varies significantly across different datasets. Specifically, the improvement is most pronounced on the Movielens-1M dataset, while it is relatively modest on the Kuairand dataset. This discrepancy is due to the substantial differences in feature quantities among the datasets. The Movielens-1M dataset has fewer features, so removing a certain amount of irrelevant features has a more significant impact and yields greater gains. A case study is given in Appendix A.6.
3.3 Data-scarcity Performance
Table 2: Data-scarcity performance of AltFS and baselines on three datasets. The metrics of best-performed methods are highlighted with bold fonts, and the sub-optimal methods are emphasized with underlined text.
Dataset Movielens-1M Aliccp Kuairand
AUC $\uparrow$ Logloss $\downarrow$ AUC $\uparrow$ Logloss $\downarrow$ AUC $\uparrow$ Logloss $\downarrow$
No Selection 0.70482 0.68468 0.55555 0.16867 0.70704 0.67961
Shallow Feature Selection
Lasso 0.71338 0.68496 0.55634 0.16812 0.71161 0.68816
GBDT 0.70131 0.66595 0.55837 0.16800 0.71398 0.67449
RF 0.70416 0.68539 0.55639 0.16846 0.71308 0.67291
XGBoost 0.71737 0.62630 0.55636 0.16865 0.71418 0.67441
Gate-based Feature Selection
AutoField 0.72051 0.62377 0.55823 0.16818 0.71256 0.67837
AdaFS 0.69554 0.65557 0.55222 0.16666 0.70571 0.67178
OptFS 0.69641 0.66066 0.55610 0.16926 0.70682 0.65146
LPFS 0.70213 0.68940 0.55768 0.16767 0.71292 0.68397
Sensitivity-based Feature Selection
Permutation 0.70700 0.69110 0.55749 0.16785 0.71370 0.67278
SFS 0.72232 0.64441 0.55570 0.16781 0.71239 0.68463
SHARK 0.70318 0.68607 0.55758 0.16798 0.71359 0.68180
AltFS 0.72372 0.62563 0.55878 0.16735 0.71463 0.66995
In this section, we discuss the performance of AltFS under data-scarcity conditions. To simulate a data-limited scenario, we reduce the training and validation sets of public datasets to 5% of their original size. Table 2 presents the results of AltFS and the baselines across three datasets. Based on Table 2, we can draw the following conclusions: 1) AltFS maintains its stable superiority even in data-scarce scenarios, almost outperforming all baselines across all metrics. This stability is due to AltFS’s ability to optimize the bridge vector with a minimal amount of data, thereby enhancing its effectiveness under limited data conditions. 2) Compared to the results in Table 1, the gap between AltFS and the No Selection approach becomes more pronounced. This indicates that feature selection has a more significant impact on the final results when data is scarce. 3) In the MovieLens-1M dataset, six baselines perform even worse than the No Selection approach. This is likely because, under data-scarce conditions, models are prone to overfitting, which prevents the agency model from providing reliable information for assessing feature importance.
3.4 Transferability Study
Table 3: Transferability study.
Model AUC $\uparrow$ Logloss $\downarrow$
No Selection AltFS No Selection AltFS
FM 0.60667 0.60725 0.16284 0.16283
DeepFM 0.61716 0.61757 0.16203 0.16180
WideDeep 0.62084 0.62212 0.16143 0.16120
DCN 0.62271 0.62322 0.16173 0.16124
In this section, we investigate the transferability of features selected by AltFS. Specifically, whether the selected features remain effective when applied to other deep recommendation backbone models. We apply the features selected by AltFS to four popular deep recommendation models: FM Rendle (2010), DeepFM Guo et al. (2017), Wide&Deep Cheng et al. (2016), and DCN Wang et al. (2017) on the Aliccp dataset. In Table 3, “No Selection” and “AltFS” represent the performance of the model with all available features and with the features selected by AltFS. We can find that: 1) All deep recommendation backbone models achieve improved prediction performance after applying the features selected by AltFS. This demonstrates that the features selected by AltFS possess strong transferability capabilities and can adapt to various deep recommender systems. 2) By comparing the performance improvements across different models after applying the AltFS-selected features, we observe that more complex models (Wide&Deep and DCN) show greater enhancements compared to FM-based models. The possible reason is that more powerful models tend to amplify the impact of the features.
3.5 Ablation Study
In this section, we compare the performance of the following variants on Aliccp to verify the effectiveness of each component: 1) w/o MM: AltFS without multiple LLMs. This variant only considers the feature sequence generated by one advanced LLM (GPT4). 2) w/o IP: AltFS without iterative prompt. This variant prompts LLM to generate feature sequences in one shot, not in an iterative manner. 3) w/o BV: AltFS without bridge vector. This variant directly derives the final feature rankings by averaging the feature sequences generated by LLMs.
Figure 4: Ablation Study.
From Figure 4, we can observe the following conclusions: 1) AltFS outperforms w/o MM. It is because the feature importance ranking from a single advanced LLM has an inherent bias, while AltFS can combine rankings from multiple LLMs to achieve a more robust and effective ranking. 2) AltFS is better than w/o IP. The possible reason is that the iterative prompt method for obtaining feature importance rankings allows the LLM to focus on a more straightforward task each time, resulting in better performance than the one-shot approach. 3) AltFS is superior to w/o BV, indicating the effectiveness of the bridge vector, which can effectively evaluate the correctness of different feature rankings and integrate them. 4) All variants outperform No Selection in all metrics, demonstrating the stability of AltFS.
3.6 Hyperparameter Analysis
Figure 5: Hyperparameter analysis on number of LLMs ($K$) and maximum masking ratio ($\beta$).
In this section, we will investigate the effect of hyperparameters on the performance of the Aliccp dataset. There are two important hyperparameters in AltFS: the number of LLMs ($K$) and the maximum masking ratio ($\beta$). Figure 5 shows the results of AUC on different hyperparameters. The full results containing both AUC and Logloss can be found in Appendix A.5.
3.6.1 Number of LLMs.
As shown in Figure 5(a), we experiment with the number of LLMs from 1 to 4. We prioritize using models with larger parameter sizes, starting with the most powerful model (GPT4). The results lead to the following conclusions: (1) A single strong LLM can provide valuable information for feature selection, with just one LLM outperforming the No selection baseline. (2) The optimal performance is achieved when the number of LLMs is set to 3. This is because the bridge vector can effectively assign weights to the existing feature importance sequences generated by different LLMs according to the current recommendation task. This results in a comprehensive feature importance ranking that integrates both semantic and task-specific aspects. (3) When the number of LLMs exceeds 3, the performance of the method stabilizes, indicating the generalization and stability of AltFS, with 3 being the most cost-effective hyperparameter choice.
3.6.2 Maximum Masking Ratio.
As illustrated in Figure 5(b), we set the maximum masking ratio from 0 to 1 and freeze other hyperparameters. We can draw the following conclusions: 1) The overall AUC trend exhibits an initial increase followed by a decrease. This is because when $\beta$ approaches 0, no features will be masked, resulting in the final feature ranking being a simple average of the sequences generated by LLMs. Conversely, when $\beta$ approaches 1, the most important features in the sequences generated by the LLMs tend to be similar at the top, while the differences among the features in the lower ranks are gradually diminished as beta increases. This weakens the effectiveness of feature selection. 2) We can also observe that the feature selection performance remains consistently exceptional when $\beta$ is between 0.2 and 0.5, highlighting the robustness and generalization capability of AltFS.
4 Related Work
Feature Selection for Recommender Systems. Features are a crucial component of machine learning algorithms, and the quality of the feature input directly impacts the model’s predictive performance Liu et al. (2019, 2023); Fan et al. (2020). Therefore, feature selection, an approach aimed at identifying predictive features from the original feature set, has become a critical part of the machine learning pipeline. Traditional feature selection methods can generally be categorized into three types: 1) Filter Methods. These methods define specific criteria for evaluating the importance of features, such as the Chi-squared test Liu and Setiono (1995) and mutual information Battiti (1994). 2) Wrapper Methods. These methods typically employ heuristic algorithms and use black-box models to assess the effectiveness of feature subsets Kohavi and John (1997). 3) Embedded Methods. This type integrates feature selection within the predictive model, evaluating features along with the model optimization process Tibshirani (1996); Friedman (2001). With the advancement of deep learning technologies, numerous related works have emerged in the field of recommender systems, leading to the development of various feature selection methods Li et al. (2024). AutoField Wang et al. (2022) was the first to introduce the use of AutoML for learning feature importance in recommender systems. It constructed an innovative controller network that learns importance weights for different feature fields. AdaFS Lin et al. (2022) expanded on AutoField by recognizing that different features may have varying importance across different samples.
AutoML in Recommender Systems. In recommender systems, challenges such as feature selection, feature crossing, and embedding dimension design face the difficulty of large candidate sets and the need for specialized knowledge to optimize them. AutoML is a technology that effectively lowers the barrier for users by automating and simplifying the design and optimization of machine learning algorithms He et al. (2021). For feature selection, methods like AutoField Wang et al. (2022), AdaFS Lin et al. (2022), and OptFS Lyu et al. (2023) have achieved significant improvements in model performance and efficiency by designing gating weights to automatically select features. In the area of feature crossing, AutoCross Luo et al. (2019) uses beam search in a tree-based space to automatically select high-level cross-features. Regarding embedding dimension search, AutoDim Zhao et al. (2021) employs a two-stage approach: it first conducts end-to-end training of the weights for each embedding dimension in a soft manner, then selects the embedding size with the highest weight for each feature field. Additionally, AutoDis Guo et al. (2021) introduces three core modules—meta-embeddings, automatic discretization, and aggregation—to learn the embedding of features.
5 Conclusion
In this paper, we introduce AltFS, an agency-light feature selection method for deep recommender systems. Specifically, we first propose a prompt iteration technique, which iteratively prompts large language models (LLMs) to obtain an initial ranking of features in the semantic space. Then, we introduce the bridge vector, which connects the semantic information of features with the current recommendation task, further fine-tuning the feature importance ranking. We conduct extensive experiments on three public datasets, and the results demonstrate the superiority of AltFS compared to other state-of-the-art methods.
6 Limitations
AltFS exhibits two potential limitations that are worth addressing in future work. First, although AltFS has optimized computational efficiency, the inference time of LLMs can still become a bottleneck when dealing with datasets containing an extremely large number of features. To mitigate this, future work could explore model quantization or distillation techniques to further enhance efficiency. Second, AltFS relies on the world knowledge encoded in LLMs, which may not fully align with the specific requirements of certain recommendation tasks. Although we have proposed the bridge vector to mitigate this issue, there is still an inevitable gap. To address this, future research could incorporate domain-specific knowledge bases or fine-tune LLMs to better adapt to specific tasks.
Appendix A Appendix
A.1 Prompt
In this section, we will detail the prompt we used in the feature importance extraction stage. The specific prompt we used for Movielens-1M is shown in Figure 6.
Figure 6: Prompt for Movielens-1M.
A.2 Dataset Description
Table 4: Dataset statistics
Dataset Movielens-1M Aliccp Kuairand
Interactions 1,000,209 85,316,519 1,436,609
Users 6,040 238,635 27,285
Items 3,706 467,298 7,551
Interaction Type Rating (1-5) Click Click
Feature Num 9 23 96
In this section, we introduce the datasets we used in our paper. The statistics of datasets are illustrated in Table 4.
  • Movielelns-1M. The Movielens-1M dataset is a highly popular dataset in the field of recommender systems, with the task of predicting whether a user will like a particular movie. In this paper, we consider interactions with a rating greater than 3 as indicative of user preference, labeling them as positive samples. This dataset consists of 9 features, which allows for the evaluation of various feature selection methods under conditions with a limited number of features. In our study, we divide the dataset into training, validation, and test sets using a 7:2:1 ratio.
  • Aliccp. The Aliccp (Alibaba Click and Conversion Prediction) dataset is derived from real-world data collected from the e-commerce platform Taobao. It contains over 80 million interaction records, encompassing more than 230,000 users and 467,000 items, making it a dataset that closely mirrors real-world scenarios. Additionally, it includes 23 features, enabling the evaluation of various feature selection methods under typical conditions. In our study, we follow the original partitioning strategy of the Aliccp dataset, using 50% of the data for training, while the remaining data is split equally between the validation and test sets.
  • Kuairand. KuaiRand is a dataset collected from the short video platform Kuaishou, based on real-world data. We have selected the “pure” version, which contains over 1 million interaction records. The primary reason for choosing this dataset is its provision of 96 features, making it well-suited for thoroughly evaluating the effectiveness of various feature selection methods in scenarios with a large number of features. We partitioned the dataset into training, validation, and test sets using a 7:2:1 ratio.
Figure 7: Hyperparameter analysis on number of LLMs ($K$) and maximum masking ratio ($\beta$) on AUC and Logloss for Aliccp dataset.
A.3 Baseline Introduction
In this section, we will provide a detailed introduction to the baselines.
A.3.1 Shallow Feature Selection Methods.
  • Lasso Tibshirani (1996). Lasso is the Least Absolute Shrinkage and Selection Operator, which effectively combines variable selection and regularization to enhance model performance.
  • GBDT Friedman (2001). Gradient Boosted Decision Trees (GBDT) is a highly effective machine learning algorithm that leverages the strengths of decision trees in combination with the boosting technique. It constructs an ensemble of decision trees, with each tree iteratively correcting the errors of its predecessors.
  • Random Forest Breiman (2001). Random Forest determines feature importance by measuring the extent to which each feature reduces impurity within a decision tree, typically using Gini impurity or entropy.
  • XGBoost Chen and Guestrin (2016). XGBoost is a highly efficient and scalable machine learning algorithm built on the principles of gradient boosting. XGBoost determines feature importance by assessing the impact of each feature on the model’s final predictive performance.
A.3.2 Gate-based Feature Selection Methods.
  • AutoField Wang et al. (2022). AutoField is the first work in the recommender systems domain to focus on feature selection within DRS (Deep Recommender Systems). It introduces an advanced controller network that trains two parameters for each feature field, representing the likelihood of selecting that feature.
  • AdaFS Lin et al. (2022). AdaFS refines feature selection further along the lines of AutoField. Recognizing that feature importance varies significantly across different samples, AdaFS focuses on generating dynamic gating weights for each sample, enabling the model to adaptively select features.
  • OptFS Lyu et al. (2023). OptFS differs from previous work by focusing on feature value selection. Through the learning of a gating network, OptFS can significantly reduce the number of feature values while simultaneously enhancing the model’s predictive performance.
  • LPFS Guo et al. (2022). LPFS critiques previous approaches that determine feature importance based on the magnitude of gating values. It introduces a smoothed-$l^{0}$ function that effectively selects features in a single stage.
A.3.3 Sensitivity-based Feature Selection Methods.
  • Permutation Fisher et al. (2019). This method assesses the importance of a feature by disrupting the values within a specific feature field and observing whether this permutation significantly impacts the model’s predictive performance.
  • SFS Wang et al. (2023). SFS assigns feature importance weights by examining the gradient of the gating vector on a small subset of data.
  • SHARK Zhang et al. (2023). SHARK uses the first-order component of the Taylor expansion as a measure of feature importance. It improves model efficiency and performance by removing less important features from the embedding table.
A.4 More details on implementation
In this section, we will introduce more details on implementation to ease reproduction.
A.4.1 Hyperparameter settings
For AltFS, we selected three LLMs: GPT-4, GPT-4o, and GPT-3.5. In the hyperparameter analysis, we also explored the performance when using four LLMs, with the additional LLM being Gemini-1.5-Pro. Our experiments were conducted on a single A100 GPU. The hyperparameter settings for other baselines are detailed in Table 5.
Table 5: Hyperparameters for baselines.
Method Hyperparameters
Shallow Feature Selection Methods
Lasso alpha=1.0
GBDT oss=logloss, learning_rate=0.1, n_estimators=100
RF n_estimators=100, criterion=gini
XGBoost booster=gbtree, gamma=0, max_depth=6
Gate-based Feature Selection Methods
AutoField update frequency: 10
AdaFs pretrained epoch: 0; update frequency: 4
OptFS gamma: 5000; epsilon: 0.1
LPFS epsilon update frequency: 100; epsilon decay: 0.9978
Sensitivity-based Feature Selection Methods
Permutation n_repeats=5
SFS num_batch_sampling: 100
SHARK None
A.4.2 Guidelines for reproduction
In this section, we primarily explain how to reproduce the results. The information is also included in the readme file in our code repository.
Feature Importance Extraction stage. You can run python llm_predict/predict.py –model=gpt-4-turbo –dataset=movielens to generate feature importance sequences by LLMs. The generating sequences will be saved automatically.
Feature Importance Refinement stage. python fs_run.py –dataset=movielens-1m –fs=altfs –retrain=false This process will run the Feature Importance Refinement stage. The terminal will output the will-trained bridge vector when the process ends.
Retraining stage. Running script by python llm_predict/llm_selection.py to get the final feature rank file. Then running python fs_run.py –dataset=movielens-1m –model=dcn –train_or_search=false –rank_path=altfs_movielens-1m –k=k –checkpoint_folder=checkpoints_for_retrain/ to retrain the recommendation model.
A.5 More results on Hyperparameter analysis
In this section, we will give the hyperparameter analysis results on both AUC and Logloss for Aliccp dataset. Since a higher AUC is better and a lower Logloss is preferable, we can observe from Figure 7 that parameter with higher AUC values correspond to lower Logloss values, indicating an inverse relationship between the two metrics. This demonstrates that our analysis of AUC in the hyperparameter analysis is equally applicable to Logloss.
Number of LLMs: Experimenting with 1 to 4 LLMs, the results reveal that a single strong LLM (e.g., GPT-4) is beneficial, outperforming the no-selection baseline. Optimal performance is achieved with 3 LLMs, as the bridge vector effectively integrates feature importance sequences from different LLMs, balancing semantic and task-specific aspects. Using more than 3 LLMs stabilizes performance, with 3 being the most cost-effective choice.
Maximum Masking Ratio: Varying the maximum masking ratio from 0.0 to 1.0 shows an initial increase in AUC followed by a decrease and an initial decrease in Logloss followed by an increase. A lower ratio (near 0) averages LLM-generated sequences, while a higher ratio (near 1) reduces the differences among features, weakening selection effectiveness. The performance remains robust between ratios of 0.2 and 0.5, indicating the stability and generalization ability of AltFS.
A.6 Case Study
Figure 8: (a) Cramér’s V matrix and (b) AUC performance on Movielens-1M.
In this section, we present a specific example to illustrate the effectiveness of LLM reasoning in the feature selection process. Figure 8(a) shows the Cramér’s V index of various features in the MovieLens-1M dataset, representing the correlation between different features. The darker the green, the stronger the correlation between the two features. We observe that “title” and “movie_id” are highly correlated, as indicated by the red box in the figure. Additionally, Figure 8(b) displays the AUC performance on the MovieLens-1M dataset for the best-performing baseline LPFS, LPFS-t (which removes the title feature based on LPFS), and AltFS. We find that removing the title feature results in a significant improvement for LPFS, indicating that the strong correlation between “title” and “movie_id” indeed hinders model training. Finally, we examine the feature importance ranking from LPFS: [“movie_id”, “user_id”, “title”, “genres”, “zip”, “gender”, “age”, “occupation”, “timestamp”]. LPFS assigns high importance to both “movie_id” and “title”, likely because the model optimization process treats these two features as roughly equally important, overlooking the issue of information redundancy between them. However, AltFS, by leveraging LLM reasoning, effectively identifies this issue. Below is an analysis extracted from the LLM response during prompt iteration regarding the title feature: “While the movie title might have some influence, it’s likely highly correlated with movie_id and thus doesn’t add significant independent information.” Consequently, a more accurate and effective feature ranking is obtained: [“user_id”, “movie_id”, “genres”, “age”, “gender”, “timestamp”, “occupation”, “zip”, “title”]. This demonstrates the effectiveness of AltFS in handling feature importance extraction.